OpenStudy (anonymous):

Solve: x - 6 = -12 A) -18 B) -6 C) 2 D) 72

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

x-6=-12 x-6 +6 = -12 +6 (add 6 to both sides) x=-6

OpenStudy (anonymous):

TANKZZZ..CAN YHU HEP ME WITH MORE

OpenStudy (anonymous):

depends. haha what's another question?